OpenStudy (anonymous):

Give the exact value of each of the following: -sin π/4

OpenStudy (anonymous):

\[\large -\sin (\frac{\pi}{4}) \implies - \frac{1}{\sqrt{2}}\] You can do rationalization here..

OpenStudy (anonymous):

Because: \[ \sin (\frac{\pi}{4}) \implies \frac{1}{\sqrt{2}}\]

OpenStudy (lgbasallote):

prove it :p

OpenStudy (anonymous):